### v1

Abstract
The recent advance in video understanding has been driven by multimodal large language models (MLLMs). But these MLLMs are good at analyzing short videos, while suffering from difficulties in understanding videos with a longer context. To address this difficulty, several agent paradigms have recently been proposed, using MLLMs as agents for retrieving extra contextual knowledge in a long video. However, most existing agents ignore the key fact that a long video is composed with multiple shots, i.e., to answer the user question from a long video, it is critical to deeply understand its relevant shots like human. Without such insight, these agents often mistakenly find redundant even noisy temporal context, restricting their capacity for long video understanding. To fill this gap, we propose VideoChat-A1, a novel long video agent paradigm. Different from the previous works, our VideoChat-A1 can deeply think with long videos, via a distinct chain-of-shot reasoning paradigm. More specifically, it can progressively select the relevant shots of user question, and look into these shots in a coarse-to-fine partition. By multi-modal reasoning along the shot chain, VideoChat-A1 can effectively mimic step-by-step human thinking process, allowing to interactively discover preferable temporal context for thoughtful understanding in long videos. Extensive experiments show that, our VideoChat-A1 achieves the state-of-the-art performance on the mainstream long video QA benchmarks, e.g., it achieves 77.0 on VideoMME and 70.1 on EgoSchema, outperforming its strong baselines (e.g., Intern2.5VL-8B
and InternVideo2.5-8B), by up to 10.8% and 6.2%. Compared to leading close-source GPT-4o and Gemini 1.5 Pro, VideoChat-A1 offers competitive accuracy, but with 7% input frames and 12% inference time on average. We will release the codes afterwards.
1 Introduction
Video understanding is an important problem in computer vision [54], which has great potential in real-world applications such as surveillance, smart phones, robotics, etc. With fast development of multimodal large language models (MLLMs), video understanding has achieved significant progress [34, 41, 44, 6, 1, 62]. By aligning video encoder with LLM, these models can flexibly interact with videos in dialogues. However, most existing MLLMs are good at understanding short videos in seconds, while hardly analyze longer videos in minutes (or hours) properly. The main challenge lies in feeding a huge amount of frames in long videos to MLLMs. To deal with this problem, several attempts have been proposed by modeling such long multimodal context [34, 41, 27], or equipping MLLMs with token compression [24, 25, 39], But their efficiency and effectiveness still need to be further improved in tackling redundant video content, thus blocking their performance on long video understanding benchmarks.
Figure 1: Motivation. Direct reasoning models such as GPT-4o [34] and Qwen2.5VL-72B [1] perform global sampling on videos and struggle to focus on key information in long videos. Agent-based methods like VideoTree [46] often suffer from incorrect or redundant sampling to generate noisy caption for wrong answer, due to the lack of deep thinking on shots in a long video. In contrast, VideoChat-A1 interactively employs shot perception and reasoning via Chain-of-Shot, which progressively looks into relevant shots through a reflective process to achieve superior performance.
Recent studies have shown that, agent-based approaches are promising for long video understanding  [43, 4, 9, 18, 28, 52, 30, 2, 60, 46]. Instead of feeding the entire long video with model adaptation, they simplify this challenge by invoking various tools to extract and retrieve relevant information from long videos. However, these approaches lack flexibility to capture complex content changes in a long video, since the retrieval is based on off-the-shelf video knowledge which is fixed after extraction. Alternatively, OpenAI o3 [35] shows a human-like thinking process on images. Via chain-of-thought reasoning, o3 can deeply understand visual content via progressive interaction with images. A similar insight has been also introduced in understanding long videos [43, 46], i.e., instead of watching a long video once, these approaches iteratively find relevant video frames for question answering. However, they ignore that, a long video consists of multiple shots. To answer the user question reliably, it is critical to deeply understand its relevant shots via multi-round reasoning. Without this consideration, these agents often mistakenly find redundant even noisy temporal context, restricting their capacity for long video understanding.
To fill this gap, we propose VideoChat-A1, a novel video agent framework that can progressively think with a long video, via an interactive chain-of-shot reasoning paradigm. Specifically, for each shot reasoning step, we leverage MLLM as core agent of thinking, and invoke tools for selecting relevant shots, dividing these shots into subshots, and reasoning user answer with these subshots. If the answer is unconfident, it means that the current subshots are insufficient for understanding user question. As a result, VideoChat-A1 will start next-step shot reasoning, in order to further understand user question with finer relevant shots. Through such a distinct chain-of-shot paradigm, our VideoChat-A1 can effectively mimic the thinking process of human, by progressively reasoning on user answer while iteratively looking into relevant video shots, as shown in Fig. 1. To show its effectiveness, we implement VideoChat-A1 on 3 strong MLLMs with 4 mainstream long video question answering benchmarks. Extensive experiments show that, our VideoChat-A1 achieves the state-of-the-art performance, e.g., it achieves 77.0 on VideoMME with subset and 70.1 on EgoSchema, outperforming its baselines such as InternVL2.5-8B and InternVideo2.5-8B, by up to 10.8% and 6.2%. Compared to GPT-4o and Gemini 1.5 Pro, VideoChat-A1 offers competitive accuracy, but with 7% input frames and 12% inference time on average.
2 Related Work
MLLM for Long Video Understanding.
Multimodal Large Models (MLLMs) have demonstrated great potential in the field of video understanding [34, 41, 58, 42, 22]. In tasks such as video question-answering, captioning, and grounding, MLLMs have significantly enhanced the understanding performance by constructing cross-modal semantic representations [25, 12, 40, 1, 6]. Despite the encouraging progress of MLLMs, they still face challenges when dealing with long-form videos that can be minutes or even hours in length and contain a substantial amount of redundant information [40]. Currently, there are two main approaches to addressing the problem of long-form videos. One is to compress visual tokens [40, 57, 47, 25, 10, 55]. For example, MA-LLM [12] merges similar tokens, and LLaMA-VID [25] compresses each frame into context and content tokens to enable longer video input. However, these methods result in a significant loss of visual information. Another approach is to extend the number of processable tokens [26, 41, 42, 51]. Models like LongVILA [51, 14, 24] have adopted strategies to increase the token capacity of LLMs (Large Language Models) for handling longer videos. However, this method introduces substantial redundant information, consumes more computational resources, and often achieves only moderate effectiveness.
Agent Based Method for Long-context Understanding.
Agent mechanisms [21, 8, 45, 16, 38, 49, 3, 4, 18, 28, 53] have been gradually introduced in video understanding tasks to enhance the semantic modeling and information acquisition capabilities for MLLMs. Currently, Agent methods for long video understanding can be mainly divided into two categories. One is extracting knowledge from long videos by invoking external tools at once and answering questions by retrieving the extracted knowledge. For example, CLIP [36] is used for retrieving key frames related to the questions [59, 50]. Memory banks [9, 12] and search engines [20, 5] are applied for extracting key information of the videos. However, this single-pass information extraction approach may lead to the omission of critical details or the inclusion of excessive redundant information, both of which interfere with the model’s response. Some methods convert the long video into long text for understanding [5, 46, 43]. The other approach involves answering questions through a round by round key information search. The multi-round video understanding pipeline emphasizes active video exploration and dynamic information acquisition by the model [46, 43, 30], and optimizes the reasoning path through a multi-round retrieval and reflection mechanism. However, a critical limitation is the lack of focus on video shots, the fundamental structural units of long-form videos, during content search. This oversight can introduce noisy information or abrupt content shifts, thereby degrading model performance.
Figure 2: Framework. VideoChat-A1 introduces a novel Chain-of-Shot Reasoning framework for long video understanding. It progressively refines video analysis through iterative stages of Shot Selection, Shot Partition, and Shot Reflection, leveraging MLLMs to dynamically discover relevant video shots and generate reliable answer.
3 Method
In this section, we introduce our VideoChat-A1 in detail. To deeply understand the user question in a long video, we design a distinct Chain-Of-Shot (CoS) reasoning paradigm, which can progressively think while discovering relevant shots via multi-round shot partition and dialogue with MLLMs.
Specifically, it consists of three key steps in each round of shot reasoning, including Shot Selection, Shot Partition, and Shot Reflection. The framework of VideoChat-A1 is shown in Fig. 2. We will explain all of them in the following.
Video Glance. To start with, it is necessary to determine if the question of the given video needs to dig into the local shots for answering, since some questions may refer to the global content throughout the entire video. In this regard, we introduce a concise video glance step for pre-judging. First, we uniformly sample 4 frames from the entire video to roughly describe the video. Then, we feed these frames along with user question and options into MLLM, justifying whether it is necessary to view the entire video to answer this question. If MLLM considers the question to be global, we uniformly sample 32 frames from the entire video, and feed them into MLLM for answering. If MLLM considers the question to be local, we start chain-of-shot reasoning.
3.1 Shot Selection
For conciseness, we describe the $i$-th shot reasoning step for illustration. Suppose that, we divide the whole video into $M$ shots at the $(i-1)$-th round,
$\mathcal{S}^{(i-1)}=\{\mathcal{S}^{(i-1)}_{1},...,\mathcal{S}^{(i-1)}_{m},...,\mathcal{S}^{(i-1)}_{M}\},$ (1)
where we treat the entire video as one single shot at the 1st round. As shown in Fig. 2, we first need to select the candidate shots from the existing ones at the $(i-1)$-th round, in order to further look into these candidates for reasoning at the $i$-th round.
Key Information Summary. To achieve this goal, we leverage MLLM to summarize key text information $\mathcal{I}^{(i)}$ to describe how to answer the user question, according to the reasoning results in the previous round,
$\mathcal{I}^{(i)}=\text{MLLM}(\mathcal{Q},~{}\mathcal{O},~{}\mathcal{V}^{(i-1)},~{}\mathcal{H}^{(i-1)}),$ (2)
where $\mathcal{Q}$ and $\mathcal{O}$ refers to the user question and answer options. $\mathcal{V}^{(i-1)}$ refers to the video frames sampled from the shots in the historical rounds, e.g., we sample 32 frames from each selected shot at round 1. In each subsequent round, we uniformly sample 8 frames from each selected shot and add them into the frame set.
Moreover, $\mathcal{H}^{(i-1)}$ refers to historical reasoning information, including key text information $\mathcal{I}^{(i-1)}$, the chosen answer $\mathcal{A}^{(i-1)}$, and the reason why to choose this answer $\mathcal{R}^{(i-1)}$ at round $(i-1)$.
Shot Selection via Retrieval. After obtaining the key text information $\mathcal{I}^{(i)}$, we leverage it as contextual guidance for shot selection. This can be effectively achieved by shot-text retrieval,
$\mathcal{C}^{(i)}=\text{CLIP}(\mathcal{I}^{(i)},~{}\mathcal{S}^{(i-1)}),$ (3)
where we leverage our finetuned LongCLIP [56] as a retriever,
and compute the cosine similarities between key information and each shot in $\mathcal{S}^{(i-1)}$. Finally, we select top $N$ shots as candidates for further investigation, $\mathcal{C}^{(i)}=\{\mathcal{C}_{1}^{(i)},...\mathcal{C}_{N}^{(i)}\}$, where $\mathcal{C}_{n}^{(i)}$ refers to the $n$-th selected shot from the shot set $\mathcal{S}^{(i-1)}$. Additionally, since there is one single shot (i.e., the whole video) at the 1st round, we directly use this shot as the candidate at this round.
3.2 Shot Partition
Figure 3: Shot Partition. Given a candidate shot at step $i$, VideoChat-A1 first applies K-Means clustering to obtain K cluster centers for finding key frames. Subsequently, subshots are partitioned based on the feature distance between each frame and its adjacent two key frames.
After obtaining the candidate shots $\mathcal{C}^{(i)}$, we next further divide each of them into subshots, in order to look deeper into it for reasoning. In this paper, we introduce a concise shot partition pipeline as shown in Fig. 3. For notation simplicity, we illustrate how to perform it on a candidate shot.
Key Frame Discovery. To discover the subshots in a candidate shot, we start by finding the key frames in the subshots. First, we uniformly sample frames from the candidate shot. To reduce temporal redundancy while maintaining computation efficiency, we sample 1 frame per second in this paper. Second, we extract the features of the sampled frames by CLIP. Third, we perform K-Means Clustering  [31] on these features. As a result, each of $K$ clusters roughly reflects a subshot. However, frames in each cluster may not be temporally adjacent, since K-Means Clustering does not take temporal order into account. To contruct subshots in a temporal order, we further find a key frame in each cluster to divide the shot. Specifically, for each cluster, we choose the frame whose feature is closest to the cluster center, as the key frame in this cluster. Finally, we organize $K$ key frames in the temporal order, $\mathcal{K}=\{\mathcal{K}(1),...,\mathcal{K}(K)\}$, for the candidate shot.
Subshot Boundary Discovery. After finding $K$ key frames, we next identify the boundary frame between two adjacent key frames to construct $K$ subshots. Specifically, we compute $\mathcal{L}_{2}$ feature distances between each frame $\mathcal{F}$ and its two adjacent key frames. Then we sum over them as a semantic deviation metric for this frame,
$d=||\mathcal{F}-\mathcal{K}(k)||_{2}+||\mathcal{F}-\mathcal{K}(k+1)||_{2},$ (4)
A higher value of $d$ indicates that this frame lies far from both adjacent key frames, implying a semantic shift between two subshots. Therefore, we select the frame with the maximum $d$ as the boundary frame between two subshots. With such a mechanism, each candidate shot is divided into $K$ subshots, $\mathcal{C}_{n}^{(i)}=\{\mathcal{S}_{n}^{(i)}(1),...,\mathcal{S}_{n}^{(i)}(K)\}$. Finally, we update the whole shot set by replacing all the candidate shots as the corresponding subshots,
$\mathcal{S}^{(i)}\leftarrow\left(\mathcal{S}^{(i-1)}\setminus\mathcal{C}^{(i)}\right)~{}\cup~{}\{\mathcal{S}_{n}^{(i)}(1),...,\mathcal{S}_{n}^{(i)}(K)\}_{n=1}^{N},$ (5)
which is used for shot selection in the next step if necessary.
3.3 Shot Reflection
Figure 4: Shot Reasoning and Shot Reflection. At each step, VideoChat-A1 performs question-answering reasoning using the relevant shots identified. It then evaluates the generated answer and historical information to reflect on the confidence level of the response. Based on the confidence score and the number of reasoning iterations, the system either proceeds to the next step for further refinement or terminates the reasoning process to output the final answer. This iterative reflection ensures reliable and contextually accurate responses.
Question Reasoning. After obtaining subshots from the candidate shots, we employ MLLMs to answer the user question based on these finer shot regions. Specifically, as shown in Fig. 4, we leverage MLLM as dialogue agent for answering question at round $i$,
$\{\mathcal{A}^{(i)},\mathcal{R}^{(i)}\}=\text{MLLM}(\mathcal{Q},~{}\mathcal{O},~{}\mathcal{V}^{(i)}),$ (6)
where $\mathcal{Q}$ and $\mathcal{O}$ refers to question and answer options. Moreover, we sample 8 frames from each subshot $\mathcal{S}_{n}^{(i)}(k)$, and combine them with $\mathcal{V}^{(i-1)}$ to form $\mathcal{V}^{(i)}$. If some sampled frames are the same as those in $\mathcal{V}^{(i-1)}$, we would merge them and resample more frames to ensure frame diversity in $\mathcal{V}^{(i)}$. Additionally, note that, we treat the whole video as the candidate shot at the 1st round without selection. If we sample frames from each subshot from this candidate, we actually look globally for question answering. Hence, we leverage key text information at the 1st round as guidance, and use CLIP to retrieve the most relevant subshot for question answering at the 1st round. Finally, given these inputs, we ask MLLM to generate answer $\mathcal{A}^{(i)}$, and the reason why to choose this answer $\mathcal{R}^{(i)}$.
Confidence Reflection. Next, MLLM would justify if the answer was reliable to decide if we need to next-round shot reasoning. Hence, we leverage MLLM again to generate the answer confidence, with extra inputs of answer and reason,
$\mathcal{Z}^{(i)}=\text{MLLM}(\mathcal{Q},~{}\mathcal{O},~{}\mathcal{V}^{(i)},\mathcal{A}^{(i)},\mathcal{R}^{(i)}).$ (7)
In this work, we set the confidence level $\mathcal{Z}^{(i)}$ from 0 to 3. If the answer confidence is higher than 2, we believe that it is a confident answer as the final output. Otherwise, we believe that it is a unreliable answer, i.e., the found subshots at this round are not sufficient to answer the question. We need to further look into them to get the answer. Hence, we start shot selection, partition, reflection in the next step. Additionally, we set a maximum number of shot reasoning steps. If VideoChat-A1 still do not get the confident answer at the max step, it will vote the major answer as the final one, according to the answers of all the rounds. Via exploit relevant shots in a coarse-to-fine manner, our VideoChat-A1 allows to leverage a chain of shots for deep thinking on long videos, and thus progressively and interactively discover the reliable shot context to answer user question.
4 Experiment
Model Egoschema LongVideoBench MLVU Video-MME
(M-avg) Short Medium Long Average
Close-Source Model
GPT-4o(0513) [34] 72.2 66.7 64.6 80.0 / 82.8 70.3 / 76.6 65.3 / 72.1 71.9 / 77.2
Gemini 1.5Pro [41] 71.1 64.0 - 81.7 / 84.5 74.3 / 81.0 67.4 / 77.4 75.0 / 81.3
Open-Source 72B-78B Model
LLaVA-OneVision-72B [19] 62.0 63.2 68.0 76.7/79.3 62.2/66.9 60.0/62.4 66.3/69.6
VideoLLaMA-2-72B [7] 63.9 - 45.6 69.8/72.0 59.9/63.0 57.6/59.0 62.4/64.7
LLAVA-Video-72B [58] 65.6 64.9 - 81.4/82.8 68.9/75.6 61.5/72.5 70.6/76.9
Qwen2.5-VL-72B [1]  77.9 - - 80.1/82.2 71.3/76.8 62.2/74.3 71.2/77.8
InternVL-2.5-78B [6] - 63.6 75.7 82.8/83.2 70.9/74.1 62.6/64.8 72.1/74.0
InternVL-3-78B [62] - 65.7 79.5 - - - 72.7 /75.7
Open-Source 7B-8B Model
ShareGPT4Video-8B [5] - 39.7 46.4 48.3/53.6 36.3/39.3 35.0/37.9 39.9/43.6
VideoChat2-7B [23] 56.7 39.3 47.9 48.3/52.8 37.0/39.4 33.2/39.2 39.5/43.8
LongVU-7B [37] 67.6 - 65.4 - - -/59.5 -/60.6
LLaVA-Video-7B [58] 57.3 58.2 70.8 - - - 63.3/69.7
Qwen2.5-VL-7B [1] 66.7 55.6 - - - - 63.3/69.0
InternVideo-2.5-8B [44] 63.9 60.6 72.8 - - - 65.1/-
InternVL-2.5-8B [6] - 60.0 68.9 - - - 64.2/66.9
InternVL-3-8B [62] - 58.8 71.4 - - - 66.3/68.9
Agent Based Model
VideoAgent [43] 54.1 - - - - - -
VideoTree (GPT-4)  [46] 61.1 - - - - - -
DrVideo [30] 61.0 - - - - - 51.7/71.7
T* [53] - - - 61.0/- 66.6/- 77.5/- 68.3/-
VideoMind [28] - - 64.4 - - 49.2/- 58.2/-
VideoMultiAgents [18] 68.0 - - - - - -
VideoRAG-72B [29] - 65.4 73.8 81.1/- 72.9/- 73.1/- 75.7/-
VideoChat-A1 (Qwen2.5-VL-7B) 70.7 64.2 71.9 78.0/81.2 73.1/77.4 64.3/70.9 71.8/76.5
VideoChat-A1 (InternVL2.5-8B) 72.1 65.2 75.1 80.8/81.4 72.3/76.8 63.7/72.6 72.3/77.0
VideoChat-A1 (InternVideo2.5-8B) 70.1 65.4 76.2 81.4/82.4 72.8/76.7 65.0/71.2 72.9/76.8
Table 1: Comparison with close-source, open-source, and agent based model on four Long Video Understanding Tasks. The Video-MME results are presented in the format “w/o subs / w/ subs".
Model Frames Inference Time EgoSchema LongVideoBench MLVU VideoMME
GPT-4o[34] 384 134.4s 72.2 66.7 64.6 71.9/77.2
Gemini-1.5-Pro[41] 568 198.8s 71.1 64.0 - 75.0/81.3
Qwen2.5-VL-72B[1] 768 122.4s 77.9 - - 71.2/77.8
Qwen2.5-VL-7B[1] 512 24.3s 66.7 55.6 - 63.3/69.7
InternVL-2.5-8B[6] 64 12.4s - 60.0 68.9 64.2/66.9
InternVideo-2.5-8B[44] 512 33.8s 63.9 60.6 72.8 65.1/-
VideoChat-A1 (Qwen2.5-VL-7B) 42.0 18.6s 70.7 64.2 71.9 71.8/76.5
VideoChat-A1 (InternVL2.5-8B) 35.9 14.7s 72.1 65.2 75.1 72.3/77.0
VideoChat-A1 (InternVideo2.5-8B) 41.3 28.4s 70.1 65.4 76.2 72.9/76.8
Table 2: Comparsion on Average Frame Number and Inference Time on Datasets. Inference Time includes the average time for video frame extraction, clustering, and the time required for the MLLM to answer all queries. Compared to closed-source or 72B/78B models, our method achieves competitive performance with shorter inference time and fewer frames. Compared to 7B/8B models, our method achieves significant performance improvements with less or comparable inference time.
4.1 Experimental Setup
Tasks & Datasets.
We test our VideoChat-A1 on four diverse long-form video question-answering benchmarks. EgoSchema [32] collects 5,032 multiple-choice question-answer pairs that span 250 hours of video content. The average video duration is 180s from an ego perspective. LongVideoBench [48] consists of 1,337 multiple-choice questions across 17 fine-grained categories average duration is 473 seconds. MLVU [61] includes 2,174 QA-pairs with an average video length of 15 minutes, ranging from 3 minutes to over 2 hours. Video-MME [11] includes 2,700 QA pairs of 900 videos. The videos’ length ranges from 3 minutes to 1 hour.
Implementation Details.
We adopt Qwen2.5-VL-7B, InternVL2.5-8B, and InternVideo2.5-8B as our baseline MLLMs.
In video glance process, we uniformly sample 4 frames. If viewing the entire video is required, we uniformly sample 32 frames for processing. In the Feature Extraction stage, we sample the video at 1 frame per second (1 FPS) for feature extraction. We use pretrained CLIP-ViT-B/32 [36] as the feature extractor, and the extracted features are used for clustering. In Shot Slection stage, we sample 16 frames from each shot as input and use a fine-tuned LongCLIP-B [56] to calculate shot-text similarity. The fine-tuning details and hyperparameters are described in the supplementary doc. We retrieve shots with similarity scores above the 0.8 threshold. In the 1st round, we select top 1 shot and sample 16 frames as initial frames at 1st round. In each subsequent round, we select top 2 shots, and divide each shot into 2 subshots. Then, we sample 8 frames from each subshot, and add them into frame set. In the Shot Partition stage, the number of clusters $K$ is set to 6 in 1st round and 2 in subsequent rounds. In the Shot Reflection stage, If the confidence score is greater than 2, the answer is considered correct; otherwise, the next iteration round continues until the confidence requirement is met or the maximum number of iterations is reached. The maximum number of iterations is set to 3. All implementations are based on Pytorch with 2 A800-80G GPUs.
4.2 Comparison with SOTA
Performance Analysis.
As shown in Table 1, we evaluate VideoChat-A1 on four mainstream benchmarks. Specifically, VideoChat-A1 significantly outperforms open-source 7-8B models, comparable to or exceeding that of large open-source or closed-source models. For example, on EgoSchema, VideoChat-A1 (InternVL2.5-8B) achieves 72.1%, surpassing InternVideo2.5-8B and VideoTree by 8.2% and 11.0%. On LongVideoBench, VideoChat-A1 (InternVideo2.5-8B) achieves 65.4%, surpassing ShareGPT4Video-8B and Gemini 1.5Pro by 25.7% and 1.4%. On MLVU, VideoChat-A1 (InternVideo2.5-8B) achieves 76.2%, outperforming GPT-4o by 11.6%. On VideoMME, VideoChat-A1 (InternVL2.5-8B) achieves 72.3% (w/o subs) and 77.0 (w/ subs), outperforming InternVL-2.5-78B by 0.2% and 3.0%.
Efficiency Analysis.
Table 2 presents a comparison of the average frame number and average inference time across various models. Compared to both closed-source models and large open-source models, VideoChat-A1 achieves substantial gains in efficiency. For instance, VideoChat-A1 (InternVL2.5-8B) processes only 35.9 frames and completes inference in 14.7 seconds on average, reducing frame usage by over 90% and inference time by more than 85% compared to GPT-4o, Gemini 1.5 Pro, and Qwen2.5-VL-72B. Despite these significant savings, our method maintains competitive performance across benchmarks. In contrast to small open-source models such as Qwen2.5-VL-7B, VideoChat-A1 not only delivers significantly better performance but also requires fewer video frames and achieves faster or comparable inference.
4.3 Ablation Study
7/8B Agent
Video
Glance
Ego LVBench MLVU VideoMME
VideoChat-A1 (Qwen2.5-VL-7B) 69.4 63.4 70.0 70.4/74.9
70.7 64.2 71.9 71.8/76.5
VideoChat-A1 (InternVL2.5-8B) 70.8 64.0 73.9 70.6/75.5
72.1 65.2 75.1 72.3/77.0
VideoChat-A1 (InternVideo2.5-8B) 69.3 64.3 74.5 71.8/74.9
70.1 65.4 76.2 72.9/76.8
Table 3: Ablation on the effect of Video Glance in VideoChat-A1.
Ablation study of Video Glance.
In this study, we conducted an ablation study to verify the impact of the video glance on the performance of the model. Table 3 shows that the model with the video glance achieves higher scores on all the tested tasks. The improvements proving the effectiveness of video glance.
Ablation on Shot Partition and Chain-of-Shot.
As shown in Table 5, we conducted an ablation study on VideoChat-A1 (Qwen2.5-VL-7B) to evaluate the impact of Shot Partition (SP) and Chain-of-Shot (CoS) across Egoschema, LongVideoBench, MLVU, and VideoMME datasets. The Egoschema and LongVideoBench are abbreviated as Ego and LVbench in the table because of the limited space. For a fair comparison, in the setup without shot partition, we evenly divided the entire video into six shots on average in the 1st round. In subsequent rounds, for the selected shots that need to be expanded, they are directly partitioned into two shots on average. Regarding the setup without Chain-of-Shot, we removed the Chain-of-Shot mechanism. Instead, the model was made to answer three rounds. In each round, the model samples additional frames from the video and incorporates them into the existing frame set from the previous round. Then we conducted a vote, selecting the answer that appeared most frequently as the final response. Individually, enabling shot partition or Chain-of-Shot boosts performance. Their combined use yields the best results. For example, on MLVU, the score rises from 66.2 (with neither) to 71.9 (both enabled), highlighting their complementary roles in enhancing long-video understanding.
SP
CoS Ego LVBench MLVU VideoMME
67.1 56.4 66.2 63.4/69.8
67.5 61.1 68.5 67.8/73.9
69.1 57.7 67.0 65.8/70.8
70.7 64.2 71.9 71.8/76.5
Table 4: Ablation of Shot Partition(SP) and Chain-of-Shot(CoS) in VideoChat-A1.
Method of
Shot Partition
Ego LVBench MLVU VideoMME
Cluster Partition 68.3 58.4 67.7 67.8/72.8
Average Partition 69.1 57.7 67.0 65.8/70.8
Shot Partition 70.7 64.2 71.9 71.8/76.5
Table 5: Ablation of different method of shot partition in VideoChat-A1 (Qwen2.5-VL-7B).
Max Rounds Ego LVBench MLVU VideoMME
1 67.3 60.1 66.2 66.8/73.4
2 70.0 62.5 69.7 69.6/76.0
3 70.7 64.2 71.9 71.8/76.5
4 70.9 64.5 72.1 72.0/76.9
Table 6: Ablation over maximum iteration rounds in VideoChat-A1 (Qwen2.5-VL-7B).
Number of
Init Shots
Ego LVBench MLVU VideoMME
3 70.1 63.9 70.8 71.2/75.7
6 70.7 64.2 71.9 71.8/76.5
9 69.3 64.0 71.8 71.5/76.2
Table 7: Ablation over the number of initialized shots in VideoChat-A1 (Qwen2.5-VL-7B).
Ablation on Shot Partition Methods.
As shown in Table 5, we evaluate the impact of three methods of shot partition across multiple benchmarks. Cluster Partition refers to the clustering approach used in VideoTree [46], where all frames within each cluster are treated as a single shot, even though they may not form a continuous video segment. Average Partition refers to dividing each shot into smaller shots by uniformly splitting it into equal parts. Our proposed Shot Partition method consistently outperforms both Cluster Partition and Average Partition across all four benchmarks. These results demonstrate that partitioning videos into temporally coherent shots based on semantic information leads to significantly better understanding of long videos. Our Shot Partition achieves best accuracies of 70.7, 64.2, 71.9, and 71.8/76.5 on EgoSchema, LongVideoBench, MLVU, and VideoMME. Compared to the better of the other two methods, our Shot Partition method achieves gains of 1.6%, 5.8%, 4.2%, and 4.0%/3.7%
Ablation Max Rounds of Reflection.
We evaluate the impact of varying the maximum number of reflection rounds across multiple benchmarks. As shown in Table 7, increasing the number of rounds consistently leads to performance improvements on all the tested dataset. For example, the accuracy on EgoSchema increases from 67.3% when using one reflection round to 70.7% with three rounds, while the accuracy on LongVideoBench improves from 60.1% to 64.2%. In the context of multi-round discussions, increasing the number of discussion rounds tends to enhance the effectiveness. After three rounds of the chain-of-shot process, the results are already quite satisfactory. As the fourth round yields only marginal improvements, we ultimately adopt the three-round setting to balance performance gains and computational cost.
Ablation on Number of Initialized Shots.
To evaluate the impact of the number of initial shots on the experimental results, the initial shot count is set to 3, 6, and 9, respectively. As shown in Table 7, the 6-shot configuration achieves best accuracies of 70.7, 64.2, 71.9, and 71.8/76.5 on EgoSchema, LongVideoBench, MLVU, and VideoMME, respectively. These correspond to gains of 1.6%, 5.8%, 4.2%, and 4.0%/3.7% over the better of the other two configurations, demonstrating that the 6-shot setting yields the best overall performance.
5 Conclusion
In this paper, we introduced VideoChat-A1, a novel agent-based framework for effectively addressing long video understanding tasks. Unlike existing approaches that often overlook the intrinsic shot-based structure of long videos, our method explicitly employs a chain-of-shot reasoning paradigm. Specifically, VideoChat-A1 progressively identifies and refines relevant video segments through three key iterative steps: Shot Selection, Shot Partition, and Shot Reflection. This iterative interaction mechanism closely mimics human cognitive processes, allowing the agent to thoughtfully reason and accurately answer user questions from long videos. Extensive experimental results demonstrate the effectiveness and efficiency of our approach, highlighting the substantial benefits of our method in long video understanding tasks.
Appendix A Prompts of VideoChat-A1.
In this section, we introduce the prompts used in VideoChat-A1, as shown in Table LABEL:tab:prompts. These prompts are carefully designed to precisely fulfill the requirements of each step in the VideoChat-A1 pipeline.
Table 8: The prompting templates used in different steps of VideoChat-A1.
Prompt Content
Prompt to decide whether to watch the whole video You are given a single-choice question, options, subtitles, and some frames of the long video. You should not only look at the textual information but also consider the input visual information, taking everything into account. If you can answer the question accurately and comprehensively based on the existing information, especially the visual information, and further watching the entire video will not significantly improve the quality of the answer, then you don’t need to watch the entire video and can answer ’No.’. However, if the existing information is not sufficient to fully answer the question, and watching the entire video may obtain information crucial for answering the question, please reply ’Yes’.
The frame tokens: {Frame tokens}
The question is:{Question}
The options is:{Options}
The subtitiles is:{Subtitiles}(If have)
Output:{Yes/No}
Prompt for generate the key information Given some frames from a long video, subtitles, a single-choice question, and options, identify the key information needed to answer the question. Focus on visual cues, context, and temporal relationships within the frames. Limit your response to 50 words.
The frame tokens: {Frame tokens}
The question is:{Question}
The options is:{Options}
The subtitiles is:{Subtitiles}(If have)
Prompt for generating the answer Given some frames from a long video, subtitles, a single-choice question, and options, select the best answer to the following question based on the video and the subtitles. Respond with only the letter (A, B, C, or D) of the correct option.
The frame tokens: {Frame tokens}
The question is:{Question}
The options is:{Options}
The subtitiles is:{Subtitiles}(If have)
The Key Info is:{Key Info}
The best answer is:
Prompt for generating the reason You are given a single-choice question, options, some frames of the long video, the key information of the question and the choice you have made. You should not only look at the textual information but also consider the input visual information, taking everything into account. Base on the provided information, your task is to explain the reason behind your choice.
The frame tokens: {Frame tokens}
The question is: {Question}
The options is:{Options}
The Key Info is:{Key Info}
Your choice is {Choice}
Prompt for confidence Evaluation You are given a single-choice question, options, some frames of the long video, your choice for the question and the reason. Your task is to evaluate whether the selected choice is correct. Criteria for Evaluation: Insufficient Information (Confidence Level: 1): There isn’t enough clear information to be sure. Partial Information (Confidence Level: 2): The decision is likely correct, but there’s some uncertainty. Sufficient Information (Confidence Level: 3): The decision is clearly supported by the given information. You need to provide the Confidence Level (a number of 1 or 2 or 3) according to the Criteria for Evaluation. Choose the confidence level that best reflects how clearly the input supports your choice. Respond with only the JSON format: ’confidence’: ’x’, where ’x’ is Confidence Level, without any additional text, explanation, or formatting.
The frame tokens: {Frame tokens}
The question is: {Question}
The options is:{Options}
Your choice is {Choice}
Your reason is {Reason}
Prompt for updating the key information in Round i Given some frames from a long video, subtitles, a single-choice question, and options, the old key information and the old choices and reasons, identify the key information needed to answer the question. Focus on visual cues, context, and temporal relationships within the frames. Limit your response to 50 words.
The frame tokens: {Frame tokens}
The question is: {Question}
Round 1 Key Info: {Key Info in Round 1}
Round 1 Choice: {Choice in Round 1}
Round 1 Reason: {Reason in Round 1}
Round i-1 Key Info: {Key Info in Round i-1}
Round i-1 Choice: {Choice in Round i-1}
Round i-1 Reason: {Reason in Round i-1}
Appendix B Visualization and Experimental Comparison of Different Models.
As shown in Fig. 5, direct reasoning models such as GPT-4o [34] and Qwen2.5VL-72B [1] perform global sampling on videos and struggle to focus on key information in long videos. Agent-based methods like VideoTree [46] often suffer from incorrect or redundant sampling to generate noisy caption for wrong answer, due to the lack of deep thinking on shots in a long video. Although VideoAgent [43] implements a dynamic frame selection process, it performs retrieval in a frame-by-frame manner, which makes it susceptible to redundant information. Moreover, VideoAgent relies heavily on captions to answer questions, resulting in the loss of substantial visual information. While CoS [13] shares the same name—Chain of Shot—with the method proposed in this work, it does not implement a multi-round dynamic shot partition. Instead, it instead performs only a single-step inference. Specifically, it uniformly samples frames from the video and groups every four adjacent frames into a single “shot.” All shots are then categorized into positive and negative shots based on their relevance to the given question. The model generates three separate answers: one from the positive shots, one from the negative shots, and one from uniformly sampled frames. These responses are aggregated to produce the final answer. In contrast, VideoChat-A1 interactively employs shot perception and reasoning via Chain-of-Shot, which progressively looks into relevant shots through a reflective process to achieve superior performance.
Table 9 presents the experimental results of our method compared with the agent-based methods shown in Fig. 5. VideoChat-A1 (InternVL2.5-8B) achieves 72.1% on Egoschema, corresponding to absolute gains of 18.0% and 11.0% over VideoAgent (54.1%) and VideoTree (61.1%), respectively. On LongVideoBench, VideoChat-A1 (InternVideo2.5-8B) attains 65.4%, improving by 12.6% over CoS (LongVA) and by 6.5% over CoS (LLaVA-Video). For MLVU (M-avg), VideoChat-A1 (InternVideo2.5-8B) reaches 76.2%, outperforming CoS (LongVA) by 17.3% and CoS (LLaVA-Video) by 4.8%. Finally, on Video-MME without subtitles, VideoChat-A1 (InternVideo2.5-8B) achieves 72.9%, with 19.4% and 8.5% improvements over CoS (LongVA) and CoS (LLaVA-Video). On Video-MME with subtitles, it achieves 76.8%, with 21.1% and 5.3% improvements over CoS (LongVA) and CoS (LLaVA-Video), respectively.
Model Egoschema LongVideoBench MLVU Video-MME
(M-avg) Short Medium Long Average
VideoAgent [43] 54.1 - - - - - -
VideoTree (GPT-4)  [46] 61.1 - - - - - -
CoS (LongVA[13] - 52.8 58.9 61.6/64.2 52.0/54.4 46.8/48.5 53.5/55.7
CoS (LLaVA-Video[13] - 58.9 71.4 77.2/80.1 62.4/69.4 53.8/65.1 64.4/71.5
VideoChat-A1 (Qwen2.5-VL-7B) 70.7 64.2 71.9 78.0/81.2 73.1/77.4 64.3/70.9 71.8/76.5
VideoChat-A1 (InternVL2.5-8B) 72.1 65.2 75.1 80.8/81.4 72.3/76.8 63.7/72.6 72.3/77.0
VideoChat-A1 (InternVideo2.5-8B) 70.1 65.4 76.2 81.4/82.4 72.8/76.7 65.0/71.2 72.9/76.8
Table 9: Comparison with other three agent-baesd models on four Long Video Understanding Tasks. The Video-MME results are presented in the format “w/o subs / w/ subs".
Figure 5: Visual comparison of different models.
Appendix C Details of the Training of LongCLIP.
To enable LongCLIP to handle long-video understanding tasks, we construct a fine-tuning dataset by sampling videos from ActivityNet [17], OpenVid-1M [33], ViTT [15], and MovieChat-Caption [40]. Specifically, we first select videos that are already segmented into multiple clips, each paired with corresponding captions. These clip-caption pairs are directly used as training samples. For the remaining videos, we filter out those shorter than 5 seconds, longer than 15 minutes, or whose captions contain fewer than 10 words. After filtering, we manually create a subset, resulting in a long-video retrieval dataset consisting of 47k samples. The average video length is 101.40 seconds, and each caption contains an average of 72.29 words. The composition of the dataset is summarized in Table 10.
Dataset Instance Num Avg Caption Length Avg Video Length
ActivityNet-Caption 2,1352 15.20 165.32
OpenVid-1M 2,2348 132.22 8.78
ViTT 3,024 22.51 280.04
MovieChat-Caption 496 132.92 432.57
Total 47,220 72.29 101.40
Table 10: Instance numbers of different datasets for training the LongCLIP.
Since LongCLIP is originally designed for image inputs, we adapt it to video by uniformly sampling 16 frames from each shot and applying a 3D convolution to extract spatiotemporal features compatible with LongCLIP’s input format. We adopt the training hyperparameters following the LongCLIP paper, with a batch size of 1024, 3 training epochs, a weight decay of 1e-2, and a learning rate of 1e-4.

### v2

@@
-
+VideoChat-A1: Thinking with Long Videos by Chain-of-Shot Reasoning
 Abstract
-The recent advance in video understanding has been driven by multimodal large language models (MLLMs).
+Recent advances in video understanding have been driven by MLLMs.
 But these MLLMs are good at analyzing short videos, while suffering from difficulties in understanding videos with a longer context.
-To address this difficulty, several agent paradigms have recently been proposed, using MLLMs as agents for retrieving extra contextual knowledge in a long video.
+To address this difficulty, several agent methods have been proposed, using MLLMs as agents for retrieving extra contextual knowledge in a long video.
 However, most existing agents ignore the key fact that a long video is composed with multiple shots, i.e., to answer the user question from a long video, it is critical to deeply understand its relevant shots like human.
@@
 More specifically, it can progressively select the relevant shots of user question, and look into these shots in a coarse-to-fine partition.
-By multi-modal reasoning along the shot chain, VideoChat-A1 can effectively mimic step-by-step human thinking process, allowing to interactively discover preferable temporal context for thoughtful understanding in long videos.
-Extensive experiments show that, our VideoChat-A1 achieves the state-of-the-art performance on the mainstream long video QA benchmarks, e.g., it achieves 77.0 on VideoMME and 70.1 on EgoSchema, outperforming its strong baselines (e.g., Intern2.5VL-8B
-and InternVideo2.5-8B), by up to 10.8% and 6.2%.
-Compared to leading close-source GPT-4o and Gemini 1.5 Pro, VideoChat-A1 offers competitive accuracy, but with 7% input frames and 12% inference time on average.
-We will release the codes afterwards.
-1 Introduction
-Video understanding is an important problem in computer vision [54], which has great potential in real-world applications such as surveillance, smart phones, robotics, etc.
-With fast development of multimodal large language models (MLLMs), video understanding has achieved significant progress [34, 41, 44, 6, 1, 62].
-By aligning video encoder with LLM, these models can flexibly interact with videos in dialogues.
-However, most existing MLLMs are good at understanding short videos in seconds, while hardly analyze longer videos in minutes (or hours) properly.
+By multi-modal reasoning along the shot chain, VideoChat-A1 can effectively mimic step-by-step human thinking process, allowing the interactive discovery of preferable temporal context for thoughtful understanding in long videos.
+Extensive experiments show that, VideoChat-A1 achieves the state-of-the-art performance on the mainstream long video QA benchmarks, e.g., it achieves 77.0 on VideoMME (w/ subs) and 70.1 on EgoSchema, outperforming its strong baselines (e.g., InternVL2.5-8B and InternVideo2.5-8B), by up to 10.1% and 6.2%.
+Compared to leading closed-source GPT-4o and Gemini 1.5 Pro, VideoChat-A1 offers competitive accuracy, but only with 7% input frames and 12% inference time on average.
+The code is available on https://github.com/SpXace/VideoChat-A1.
+Introduction
+Video understanding is an important problem in computer vision (You et al.
+2024).
+With fast development of MLLMs, video understanding has achieved significant progress (OpenAI 2024; Wang et al.
+2025).
+However, most existing MLLMs are good at understanding short videos in seconds, while hardly analyzing longer videos in minutes (or hours) properly.
 The main challenge lies in feeding a huge amount of frames in long videos to MLLMs.
-To deal with this problem, several attempts have been proposed by modeling such long multimodal context [34, 41, 27], or equipping MLLMs with token compression [24, 25, 39], But their efficiency and effectiveness still need to be further improved in tackling redundant video content, thus blocking their performance on long video understanding benchmarks.
+To deal with this problem, several attempts have been proposed by modeling such long multimodal context (OpenAI 2024; Liu et al.
+2025a), or equipping MLLMs with token compression (Li et al.
+2025a; Shu et al.
+2024), But their efficiency and effectiveness still need to be further improved in tackling redundant video content, thus blocking their performance on long video understanding benchmarks.
 Figure 1: Motivation.
-Direct reasoning models such as GPT-4o [34] and Qwen2.5VL-72B [1] perform global sampling on videos and struggle to focus on key information in long videos.
-Agent-based methods like VideoTree [46] often suffer from incorrect or redundant sampling to generate noisy caption for wrong answer, due to the lack of deep thinking on shots in a long video.
-In contrast, VideoChat-A1 interactively employs shot perception and reasoning via Chain-of-Shot, which progressively looks into relevant shots through a reflective process to achieve superior performance.
-Recent studies have shown that, agent-based approaches are promising for long video understanding  [43, 4, 9, 18, 28, 52, 30, 2, 60, 46].
+Direct reasoning models such as GPT-4o (OpenAI 2024) perform global sampling and struggle to focus on key information.
+Agent-based methods like VideoTree (Wang et al.
+2024c) often suffer from incorrect or redundant sampling that leads to noisy captions.
+In contrast, VideoChat-A1 interactively employs shot perception and reasoning via Chain-of-Shot, which progressively looks into relevant shots through a reflective process for better performance.
+Recent studies have shown that, agent-based approaches are promising for long video understanding  (Wang et al.
+2024a; Chen et al.
+2025; Kugo et al.
+2025; Zhi et al.
+2025; Wang et al.
+2024c).
 Instead of feeding the entire long video with model adaptation, they simplify this challenge by invoking various tools to extract and retrieve relevant information from long videos.
 However, these approaches lack flexibility to capture complex content changes in a long video, since the retrieval is based on off-the-shelf video knowledge which is fixed after extraction.
-Alternatively, OpenAI o3 [35] shows a human-like thinking process on images.
+Alternatively, OpenAI o3 (OpenAI et al.
+2025) shows a human-like thinking process on images.
 Via chain-of-thought reasoning, o3 can deeply understand visual content via progressive interaction with images.
-A similar insight has been also introduced in understanding long videos [43, 46], i.e., instead of watching a long video once, these approaches iteratively find relevant video frames for question answering.
 However, they ignore that, a long video consists of multiple shots.
@@
 To fill this gap, we propose VideoChat-A1, a novel video agent framework that can progressively think with a long video, via an interactive chain-of-shot reasoning paradigm.
-Specifically, for each shot reasoning step, we leverage MLLM as core agent of thinking, and invoke tools for selecting relevant shots, dividing these shots into subshots, and reasoning user answer with these subshots.
+Specifically, for each shot reasoning step, we leverage MLLM as core agent of thinking, and invoke tools for selecting relevant shots, dividing these shots into subshots, and reasoning answer with subshots.
 If the answer is unconfident, it means that the current subshots are insufficient for understanding user question.
-As a result, VideoChat-A1 will start next-step shot reasoning, in order to further understand user question with finer relevant shots.
+As a result, VideoChat-A1 will start next-step shot reasoning to further understand user question with finer relevant shots.
 Through such a distinct chain-of-shot paradigm, our VideoChat-A1 can effectively mimic the thinking process of human, by progressively reasoning on user answer while iteratively looking into relevant video shots, as shown in Fig.
 1.
-To show its effectiveness, we implement VideoChat-A1 on 3 strong MLLMs with 4 mainstream long video question answering benchmarks.
-Extensive experiments show that, our VideoChat-A1 achieves the state-of-the-art performance, e.g., it achieves 77.0 on VideoMME with subset and 70.1 on EgoSchema, outperforming its baselines such as InternVL2.5-8B and InternVideo2.5-8B, by up to 10.8% and 6.2%.
+Extensive experiments show that, our VideoChat-A1 achieves the state-of-the-art performance on 4 mainstream benchmarks, e.g., it achieves 77.0 on VideoMME with subset and 70.1 on EgoSchema, outperforming its baselines such as InternVL2.5-8B and InternVideo2.5-8B, by up to 10.1% and 6.2%.
 Compared to GPT-4o and Gemini 1.5 Pro, VideoChat-A1 offers competitive accuracy, but with 7% input frames and 12% inference time on average.
-2 Related Work
+Related Work
 MLLM for Long Video Understanding.
-Multimodal Large Models (MLLMs) have demonstrated great potential in the field of video understanding [34, 41, 58, 42, 22].
-In tasks such as video question-answering, captioning, and grounding, MLLMs have significantly enhanced the understanding performance by constructing cross-modal semantic representations [25, 12, 40, 1, 6].
-Despite the encouraging progress of MLLMs, they still face challenges when dealing with long-form videos that can be minutes or even hours in length and contain a substantial amount of redundant information [40].
-Currently, there are two main approaches to addressing the problem of long-form videos.
-One is to compress visual tokens [40, 57, 47, 25, 10, 55].
-For example, MA-LLM [12] merges similar tokens, and LLaMA-VID [25] compresses each frame into context and content tokens to enable longer video input.
+MLLMs have demonstrated great potential in the field of video understanding (OpenAI 2024; Li et al.
+2023; Chen et al.
+2022).
+However, they still face challenges when dealing with long videos that can be minutes or hours in length  (Song et al.
+2024).
+Currently, there are two main approaches to addressing this problem.
+One is to compress visual tokens (Song et al.
+2024; Zeng et al.
+2024).
+For example, MA-LLM (He et al.
+2024) merges similar tokens, and LLaMA-VID (Li et al.
+2025b) compresses each frame into context and content tokens to enable longer video input.
 However, these methods result in a significant loss of visual information.
-Another approach is to extend the number of processable tokens [26, 41, 42, 51].
-Models like LongVILA [51, 14, 24] have adopted strategies to increase the token capacity of LLMs (Large Language Models) for handling longer videos.
-However, this method introduces substantial redundant information, consumes more computational resources, and often achieves only moderate effectiveness.
+Another approach is to extend the number of processable tokens (Liu et al.
+2024; Wang et al.
+2024b).
+Models like LongVILA (Xue et al.
+2024; Li et al.
+2025a) have adopted strategies to increase the token capacity for handling longer videos.
+However, this method introduces redundant information, consumes more computational resources, and often achieves only moderate effectiveness.
 Agent Based Method for Long-context Understanding.
-Agent mechanisms [21, 8, 45, 16, 38, 49, 3, 4, 18, 28, 53] have been gradually introduced in video understanding tasks to enhance the semantic modeling and information acquisition capabilities for MLLMs.
+Agent mechanisms (Chen et al.
+2025; Kugo et al.
+2025; Ye et al.
+2025) have been introduced in video understanding tasks.
 Currently, Agent methods for long video understanding can be mainly divided into two categories.
 One is extracting knowledge from long videos by invoking external tools at once and answering questions by retrieving the extracted knowledge.
-For example, CLIP [36] is used for retrieving key frames related to the questions [59, 50].
-Memory banks [9, 12] and search engines [20, 5] are applied for extracting key information of the videos.
-However, this single-pass information extraction approach may lead to the omission of critical details or the inclusion of excessive redundant information, both of which interfere with the model’s response.
-Some methods convert the long video into long text for understanding [5, 46, 43].
+For example, CLIP (Radford et al.
+2021) is used for retrieving key frames related to the questions (Zhao et al.
+2024; Xie et al.
+2023).
+Memory banks (Fan et al.
+2024; He et al.
+2024) and search engines (Li et al.
+2024b; Chen et al.
+2024a) are applied for extracting key information of the videos.
+However, this single-pass information extraction often misses critical details or introduces redundant information, which interferes with the model’s response.
 The other approach involves answering questions through a round by round key information search.
-The multi-round video understanding pipeline emphasizes active video exploration and dynamic information acquisition by the model [46, 43, 30], and optimizes the reasoning path through a multi-round retrieval and reflection mechanism.
-However, a critical limitation is the lack of focus on video shots, the fundamental structural units of long-form videos, during content search.
-This oversight can introduce noisy information or abrupt content shifts, thereby degrading model performance.
+The multi-round pipeline emphasizes active video exploration and dynamic information acquisition (Wang et al.
+2024c, a; Ma et al.
+2024), and optimizes the reasoning path through retrieval and reflection mechanism.
+However, a critical limitation is that most methods either ignore video shots—the fundamental structural units of long-form videos—or segment shots at once without further fine-grained partitioning, which can both miss critical query-relevant details within each shot and introduce redundant or noisy information.
+For instance, VideoINSTA (Liao et al.
+2024) proposes an event-based spatial-temporal reasoning framework, but its spatial and temporal information extraction processes are relatively independent and lack query-driven interactive iterative refinement.
 Figure 2: Framework.
@@
 It progressively refines video analysis through iterative stages of Shot Selection, Shot Partition, and Shot Reflection, leveraging MLLMs to dynamically discover relevant video shots and generate reliable answer.
-3 Method
-In this section, we introduce our VideoChat-A1 in detail.
-To deeply understand the user question in a long video, we design a distinct Chain-Of-Shot (CoS) reasoning paradigm, which can progressively think while discovering relevant shots via multi-round shot partition and dialogue with MLLMs.
+Method
+In this section, we introduce our VideoChat-A1.
+To deeply understand the user question in a long video, we design a distinct Chain-of-Shot (CoS) reasoning paradigm, which can progressively think while discovering relevant shots via multi-round shot partition and dialogues.
 Specifically, it consists of three key steps in each round of shot reasoning, including Shot Selection, Shot Partition, and Shot Reflection.
@@
 2.
-We will explain all of them in the following.
 Video Glance.
@@
 If MLLM considers the question to be local, we start chain-of-shot reasoning.
-3.1 Shot Selection
-For conciseness, we describe the $i$-th shot reasoning step for illustration. Suppose that, we divide the whole video into $M$ shots at the $(i-1)$-th round,
+Shot Selection
+For conciseness, we describe the $i$-th shot reasoning step for illustration. Suppose that, we divide the whole video into $M$ shots at the (i1)(i-1)-th round,
 $\mathcal{S}^{(i-1)}=\{\mathcal{S}^{(i-1)}_{1},...,\mathcal{S}^{(i-1)}_{m},...,\mathcal{S}^{(i-1)}_{M}\}$, (1)
-where we treat the entire video as one single shot at the 1st round.
-As shown in Fig.
-2, we first need to select the candidate shots from the existing ones at the $(i-1)$-th round, in order to further look into these candidates for reasoning at the $i$-th round.
+As shown in Fig. 2, we first select the candidate shots from the existing ones at the (i1)(i-1)-th round, in order to further look into these candidates for reasoning at the $i$-th round.
 Key Information Summary.
@@
 where $\mathcal{Q}$ and $\mathcal{O}$ refers to the user question and answer options.
-$\mathcal{V}^{(i-1)}$ refers to the video frames sampled from the shots in the historical rounds, e.g., we sample 32 frames from each selected shot at round 1.
-In each subsequent round, we uniformly sample 8 frames from each selected shot and add them into the frame set.
-Moreover, $\mathcal{H}^{(i-1)}$ refers to historical reasoning information, including key text information $\mathcal{I}^{(i-1)}$, the chosen answer $\mathcal{A}^{(i-1)}$, and the reason why to choose this answer $\mathcal{R}^{(i-1)}$ at round $(i-1)$.
+$\mathcal{V}^{(i-1)}$ refers to the video frames sampled from the shots in the historical rounds.
+Moreover, $\mathcal{H}^{(i-1)}$ refers to historical reasoning information, including key text information $\mathcal{I}^{(i-1)}$, the chosen answer $\mathcal{A}^{(i-1)}$, and the reason why to choose this answer $\mathcal{R}^{(i-1)}$ at round (i1)(i-1).
 Shot Selection via Retrieval.
@@
 $\mathcal{C}^{(i)}=\text{CLIP}(\mathcal{I}^{(i)},~\mathcal{S}^{(i-1)}),$ (3)
-where we leverage our finetuned LongCLIP [56] as a retriever,
-and compute the cosine similarities between key information and each shot in $\mathcal{S}^{(i-1)}$.
+where we leverage our finetuned LongCLIP (Zhang et al.
+2024a) as a retriever, and compute the cosine similarities between key information and each shot in $\mathcal{S}^{(i-1)}$.
 Finally, we select top $N$ shots as candidates for further investigation, $\mathcal{C}^{(i)}=\{\mathcal{C}_{1}^{(i)},...\mathcal{C}_{N}^{(i)}\}$, where $\mathcal{C}_{n}^{(i)}$ refers to the $n$-th selected shot from the shot set $\mathcal{S}^{(i-1)}$.
 Additionally, since there is one single shot (i.e., the whole video) at the 1st round, we directly use this shot as the candidate at this round.
-3.2 Shot Partition
+Shot Partition
 Figure 3: Shot Partition.
-Given a candidate shot at step $i$, VideoChat-A1 first applies K-Means clustering to obtain K cluster centers for finding key frames.
-Subsequently, subshots are partitioned based on the feature distance between each frame and its adjacent two key frames.
-After obtaining the candidate shots $\mathcal{C}^{(i)}$, we next further divide each of them into subshots, in order to look deeper into it for reasoning.
+Given a candidate shot at step $i$, VideoChat-A1 first applies K-Means to obtain K cluster centers for finding key frames.
+Subsequently, subshots are partitioned based on the feature distance and its adjacent two key frames.
+After obtaining the candidate shots $\mathcal{C}^{(i)}$, we next further divide each of them into subshots, to look deeper into it for reasoning.
 In this paper, we introduce a concise shot partition pipeline as shown in Fig.
@@
 Key Frame Discovery.
-To discover the subshots in a candidate shot, we start by finding the key frames in the subshots.
+To discover the subshots in a candidate shot, we start by finding key frames in the subshots.
 First, we uniformly sample frames from the candidate shot.
-To reduce temporal redundancy while maintaining computation efficiency, we sample 1 frame per second in this paper.
+To reduce temporal redundancy while maintaining computation efficiency, we sample 1 fps in this paper.
 Second, we extract the features of the sampled frames by CLIP.
-Third, we perform K-Means Clustering  [31] on these features.
+Third, we perform K-Means (MacQueen 1967) on these features.
 As a result, each of $K$ clusters roughly reflects a subshot.
+The choice of $K$ is discussed in detail in the ablation study section.
 However, frames in each cluster may not be temporally adjacent, since K-Means Clustering does not take temporal order into account.
-To contruct subshots in a temporal order, we further find a key frame in each cluster to divide the shot.
+To construct subshots in temporal order, we further find a key frame in each cluster to divide the shot.
 Specifically, for each cluster, we choose the frame whose feature is closest to the cluster center, as the key frame in this cluster.
@@
 Then we sum over them as a semantic deviation metric for this frame,
-$d=||\mathcal{F}-\mathcal{K}(k)||_{2}+||\mathcal{F}-\mathcal{K}(k+1)||_{2},$ (4)
+[…]
 A higher value of $d$ indicates that this frame lies far from both adjacent key frames, implying a semantic shift between two subshots.
@@
 which is used for shot selection in the next step if necessary.
-3.3 Shot Reflection
-Figure 4: Shot Reasoning and Shot Reflection.
-At each step, VideoChat-A1 performs question-answering reasoning using the relevant shots identified.
-It then evaluates the generated answer and historical information to reflect on the confidence level of the response.
-Based on the confidence score and the number of reasoning iterations, the system either proceeds to the next step for further refinement or terminates the reasoning process to output the final answer.
-This iterative reflection ensures reliable and contextually accurate responses.
+Shot Reflection
 Question Reasoning.
@@
 Specifically, as shown in Fig.
-4, we leverage MLLM as dialogue agent for answering question at round $i$,
+4, we leverage MLLM as a dialogue agent to answer the question at round $i$,
 $\{\mathcal{A}^{(i)},\mathcal{R}^{(i)}\}=\text{MLLM}(\mathcal{Q},~\mathcal{O},~\mathcal{V}^{(i)}),$ (6)
-where $\mathcal{Q}$ and $\mathcal{O}$ refers to question and answer options.
-Moreover, we sample 8 frames from each subshot $\mathcal{S}_{n}^{(i)}(k)$, and combine them with $\mathcal{V}^{(i-1)}$ to form $\mathcal{V}^{(i)}$.
-If some sampled frames are the same as those in $\mathcal{V}^{(i-1)}$, we would merge them and resample more frames to ensure frame diversity in $\mathcal{V}^{(i)}$.
-Additionally, note that, we treat the whole video as the candidate shot at the 1st round without selection.
-If we sample frames from each subshot from this candidate, we actually look globally for question answering.
-Hence, we leverage key text information at the 1st round as guidance, and use CLIP to retrieve the most relevant subshot for question answering at the 1st round.
-Finally, given these inputs, we ask MLLM to generate answer $\mathcal{A}^{(i)}$, and the reason why to choose this answer $\mathcal{R}^{(i)}$.
+where $\mathcal{Q}$ and $\mathcal{O}$ refer to question and answer options.
+We sample $\mathcal{T}$ frames from each subshot $\mathcal{S}_{n}^{(i)}(k)$, where $\mathcal{T}$ is the number of sampled frames, and combine them with $\mathcal{V}^{(i-1)}$ to form $\mathcal{V}^{(i)}$.
+Finally, given these inputs, MLLM generates answer $\mathcal{A}^{(i)}$, and the reason why to choose this answer $\mathcal{R}^{(i)}$.
 Confidence Reflection.
@@
 $\mathcal{Z}^{(i)}=\text{MLLM}(\mathcal{Q},~\mathcal{O},~\mathcal{V}^{(i)},\mathcal{A}^{(i)},\mathcal{R}^{(i)}).$ (7)
-In this work, we set the confidence level $\mathcal{Z}^{(i)}$ from 0 to 3.
-If the answer confidence is higher than 2, we believe that it is a confident answer as the final output.
-Otherwise, we believe that it is a unreliable answer, i.e., the found subshots at this round are not sufficient to answer the question.
-We need to further look into them to get the answer.
-Hence, we start shot selection, partition, reflection in the next step.
-Additionally, we set a maximum number of shot reasoning steps.
-If VideoChat-A1 still do not get the confident answer at the max step, it will vote the major answer as the final one, according to the answers of all the rounds.
+We set the confidence level $\mathcal{Z}^{(i)}$ from 0 to 3.
+If the confidence is higher than 2, we believe that it is a reliable answer as the final output.
+Otherwise, we believe that it is an unreliable answer and conduct further investigation.
+Hence, we start shot selection, partition, reflection in the next round.
+Additionally, we set a maximum number of shot reasoning rounds.
+If VideoChat-A1 still does not get the confident answer at the max step, it will vote the major answer as the final one, according to the answers of all the rounds.
 Via exploit relevant shots in a coarse-to-fine manner, our VideoChat-A1 allows to leverage a chain of shots for deep thinking on long videos, and thus progressively and interactively discover the reliable shot context to answer user question.
-4 Experiment
 Model EgoSchema LongVideoBench MLVU Video-MME
 (M-avg) Short Medium Long Average
-Close-Source Model
-GPT-4o(0513) [34] 72.2 66.7 64.6 80.0 / 82.8 70.3 / 76.6 65.3 / 72.1 71.9 / 77.2
-Gemini 1.5Pro [41] 71.1 64.0 - 81.7 / 84.5 74.3 / 81.0 67.4 / 77.4 75.0 / 81.3
+Closed-Source Model
+GPT-4o(0513) (OpenAI 2024) 72.2 66.7 64.6 80.0 / 82.8 70.3 / 76.6 65.3 / 72.1 71.9 / 77.2
+Gemini 1.5 Pro (Reid et al. 2024) 71.1 64.0 - 81.7 / 84.5 74.3 / 81.0 67.4 / 77.4 75.0 / 81.3
 Open-Source 72B-78B Model
-LLaVA-OneVision-72B [19] 62.0 63.2 68.0 76.7/79.3 62.2/66.9 60.0/62.4 66.3/69.6
-VideoLLaMA-2-72B [7] 63.9 - 45.6 69.8/72.0 59.9/63.0 57.6/59.0 62.4/64.7
-LLAVA-Video-72B [58] 65.6 64.9 - 81.4/82.8 68.9/75.6 61.5/72.5 70.6/76.9
-Qwen2.5-VL-72B [1]  77.9 - - 80.1/82.2 71.3/76.8 62.2/74.3 71.2/77.8
-InternVL-2.5-78B [6] - 63.6 75.7 82.8/83.2 70.9/74.1 62.6/64.8 72.1/74.0
-InternVL-3-78B [62] - 65.7 79.5 - - - 72.7 /75.7
+LLaVA-OneVision-72B (Li et al. 2024a) 62.0 63.2 68.0 76.7/79.3 62.2/66.9 60.0/62.4 66.3/69.6
+VideoLLaMA-2-72B (Cheng et al. 2024) 63.9 - 45.6 69.8/72.0 59.9/63.0 57.6/59.0 62.4/64.7
+LLAVA-Video-72B (Zhang et al. 2024b) 65.6 64.9 - 81.4/82.8 68.9/75.6 61.5/72.5 70.6/76.9
+Qwen2.5-VL-72B (Bai et al. 2025)  77.9 - - 80.1/82.2 71.3/76.8 62.2/74.3 71.2/77.8
+InternVL-2.5-78B (Chen et al. 2024b) - 63.6 75.7 82.8/83.2 70.9/74.1 62.6/64.8 72.1/74.0
+InternVL-3-78B (Zhu et al. 2025) - 65.7 79.5 - - - 72.7 /75.7
 Open-Source 7B-8B Model
-ShareGPT4Video-8B [5] - 39.7 46.4 48.3/53.6 36.3/39.3 35.0/37.9 39.9/43.6
-VideoChat2-7B [23] 56.7 39.3 47.9 48.3/52.8 37.0/39.4 33.2/39.2 39.5/43.8
-LongVU-7B [37] 67.6 - 65.4 - - -/59.5 -/60.6
-LLaVA-Video-7B [58] 57.3 58.2 70.8 - - - 63.3/69.7
-Qwen2.5-VL-7B [1] 66.7 55.6 - - - - 63.3/69.0
-InternVideo-2.5-8B [44] 63.9 60.6 72.8 - - - 65.1/-
-InternVL-2.5-8B [6] - 60.0 68.9 - - - 64.2/66.9
-InternVL-3-8B [62] - 58.8 71.4 - - - 66.3/68.9
+ShareGPT4Video-8B (Chen et al. 2024a) - 39.7 46.4 48.3/53.6 36.3/39.3 35.0/37.9 39.9/43.6
+VideoChat2-7B (Li et al. 2024c) 56.7 39.3 47.9 48.3/52.8 37.0/39.4 33.2/39.2 39.5/43.8
+LongVU-7B (Shen et al. 2024) 67.6 - 65.4 - - -/59.5 -/60.6
+LLaVA-Video-7B (Zhang et al. 2024b) 57.3 58.2 70.8 - - - 63.3/69.7
+Qwen2.5-VL-7B (Bai et al. 2025) 65.0 56.0 - - - - 65.1/71.6
+InternVideo-2.5-8B (Wang et al. 2025) 63.9 60.6 72.8 - - - 65.1/-
+InternVL-2.5-8B (Chen et al. 2024b) - 60.0 68.9 - - - 64.2/66.9
+InternVL-3-8B (Zhu et al. 2025) - 58.8 71.4 - - - 66.3/68.9
 Agent Based Model
-VideoAgent [43] 54.1 - - - - - -
-VideoTree (GPT-4)  [46] 61.1 - - - - - -
-DrVideo [30] 61.0 - - - - - 51.7/71.7
-T* [53] - - - 61.0/- 66.6/- 77.5/- 68.3/-
-VideoMind [28] - - 64.4 - - 49.2/- 58.2/-
-VideoMultiAgents [18] 68.0 - - - - - -
-VideoRAG-72B [29] - 65.4 73.8 81.1/- 72.9/- 73.1/- 75.7/-
+VideoAgent (Wang et al. 2024a) 54.1 - - - - - -
+VideoTree (GPT-4)  (Wang et al. 2024c) 61.1 - - - - - -
+DrVideo (Ma et al. 2024) 61.0 - - - - - 51.7/71.7
+T* (Ye et al. 2025) - - - 61.0/- 66.6/- 77.5/- 68.3/-
+VideoMind (Liu et al. 2025b) - - 64.4 - - 49.2/- 58.2/-
+VideoMultiAgents (Kugo et al. 2025) 68.0 - - - - - -
+VideoRAG-72B (Luo et al. 2024) - 65.4 73.8 81.1/- 72.9/- 73.1/- 75.7/-
 VideoChat-A1 (Qwen2.5-VL-7B) 70.7 64.2 71.9 78.0/81.2 73.1/77.4 64.3/70.9 71.8/76.5
@@
 VideoChat-A1 (InternVideo2.5-8B) 70.1 65.4 76.2 81.4/82.4 72.8/76.7 65.0/71.2 72.9/76.8
-Table 1: Comparison with close-source, open-source, and agent based model on four Long Video Understanding Tasks. The Video-MME results are presented in the format “w/o subs / w/ subs".
+Table 1: Comparison with closed-source, open-source, and agent based model on four Long Video Understanding Tasks. The Video-MME results are presented in the format “w/o subs / w/ subs".
 Model Frames Inference Time EgoSchema LongVideoBench MLVU VideoMME
-GPT-4o[34] 384 134.4s 72.2 66.7 64.6 71.9/77.2
-Gemini-1.5-Pro[41] 568 198.8s 71.1 64.0 - 75.0/81.3
-Qwen2.5-VL-72B[1] 768 122.4s 77.9 - - 71.2/77.8
-Qwen2.5-VL-7B[1] 512 24.3s 66.7 55.6 - 63.3/69.7
-InternVL-2.5-8B[6] 64 12.4s - 60.0 68.9 64.2/66.9
-InternVideo-2.5-8B[44] 512 33.8s 63.9 60.6 72.8 65.1/-
+GPT-4o(OpenAI 2024) 384 134.4s 72.2 66.7 64.6 71.9/77.2
+Gemini-1.5-Pro(Reid et al. 2024) 568 198.8s 71.1 64.0 - 75.0/81.3
+Qwen2.5-VL-72B(Bai et al. 2025) 768 122.4s 77.9 - - 71.2/77.8
+Qwen2.5-VL-7B(Bai et al. 2025) 512 24.3s 66.7 55.6 - 63.3/69.7
+InternVL-2.5-8B(Chen et al. 2024b) 64 12.4s - 60.0 68.9 64.2/66.9
+InternVideo-2.5-8B(Wang et al. 2025) 512 33.8s 63.9 60.6 72.8 65.1/-
 VideoChat-A1 (Qwen2.5-VL-7B) 42.0 18.6s 70.7 64.2 71.9 71.8/76.5
@@
 VideoChat-A1 (InternVideo2.5-8B) 41.3 28.4s 70.1 65.4 76.2 72.9/76.8
-Table 2: Comparsion on Average Frame Number and Inference Time on Datasets.
+Table 2: Comparison on Average Frame Number and Inference Time on Datasets.
 Inference Time includes the average time for video frame extraction, clustering, and the time required for the MLLM to answer all queries.
-Compared to closed-source or 72B/78B models, our method achieves competitive performance with shorter inference time and fewer frames.
-Compared to 7B/8B models, our method achieves significant performance improvements with less or comparable inference time.
-4.1 Experimental Setup
-Tasks & Datasets.
-We test our VideoChat-A1 on four diverse long-form video question-answering benchmarks.
-EgoSchema [32] collects 5,032 multiple-choice question-answer pairs that span 250 hours of video content.
-The average video duration is 180s from an ego perspective.
-LongVideoBench [48] consists of 1,337 multiple-choice questions across 17 fine-grained categories average duration is 473 seconds.
-MLVU [61] includes 2,174 QA-pairs with an average video length of 15 minutes, ranging from 3 minutes to over 2 hours.
-Video-MME [11] includes 2,700 QA pairs of 900 videos.
-The videos’ length ranges from 3 minutes to 1 hour.
-Implementation Details.
-We adopt Qwen2.5-VL-7B, InternVL2.5-8B, and InternVideo2.5-8B as our baseline MLLMs.
-In video glance process, we uniformly sample 4 frames.
-If viewing the entire video is required, we uniformly sample 32 frames for processing.
-In the Feature Extraction stage, we sample the video at 1 frame per second (1 FPS) for feature extraction.
-We use pretrained CLIP-ViT-B/32 [36] as the feature extractor, and the extracted features are used for clustering.
-In Shot Slection stage, we sample 16 frames from each shot as input and use a fine-tuned LongCLIP-B [56] to calculate shot-text similarity.
-The fine-tuning details and hyperparameters are described in the supplementary doc.
+Experiment
+Experimental Setup
+We evaluate VideoChat-A1 on four long-video question-answering benchmarks: EgoSchema (Mangalam et al.
+2023), LongVideoBench (Wu et al.
+2024), MLVU (Zhou et al.
+2024), and Video-MME (Fu et al.
+2024).
+As baselines, we use Qwen2.5-VL-7B, InternVL2.5-8B, and InternVideo2.5-8B.
+In the Feature Extraction stage, we sample the video at 1 fps and use pretrained CLIP-ViT-B/32 (Radford et al.
+2021) as the feature extractor.
+In Shot Selection, we sample 16 frames from each shot and use a fine-tuned LongCLIP-B (Zhang et al.
+2024a) to calculate shot-text similarity.
+The details are described in the supplementary documentation.
 We retrieve shots with similarity scores above the 0.8 threshold.
-In the 1st round, we select top 1 shot and sample 16 frames as initial frames at 1st round.
-In each subsequent round, we select top 2 shots, and divide each shot into 2 subshots.
-Then, we sample 8 frames from each subshot, and add them into frame set.
+In the first round, we treat the whole video as the candidate shot without selection, and sample $\mathcal{T}=16$ frames as initial frames.
+In each subsequent round, we select top 2 shots, divide each shot into 2 subshots, and sample $\mathcal{T}=8$ frames from each subshot, and add them into frame set.
 In the Shot Partition stage, the number of clusters $K$ is set to 6 in 1st round and 2 in subsequent rounds.
-In the Shot Reflection stage, If the confidence score is greater than 2, the answer is considered correct; otherwise, the next iteration round continues until the confidence requirement is met or the maximum number of iterations is reached.
-The maximum number of iterations is set to 3.
-All implementations are based on Pytorch with 2 A800-80G GPUs.
-4.2 Comparison with SOTA
-Performance Analysis.
+The maximum number of iterations is 3.
+All experiments are conducted on 2 A800-80GB GPUs.
+Figure 4: Shot Reasoning and Shot Reflection.
+At each round, VideoChat-A1 performs question-answering reasoning using the relevant shots identified.
+It then evaluates the confidence level of the response.
+Then, the system either proceeds to the next step for further refinement or terminates the reasoning process to output the final answer.
+Comparison with SOTA
+Performance and Efficiency Analysis.
 As shown in Table 1, we evaluate VideoChat-A1 on four mainstream benchmarks.
 Specifically, VideoChat-A1 significantly outperforms open-source 7-8B models, comparable to or exceeding that of large open-source or closed-source models.
-For example, on EgoSchema, VideoChat-A1 (InternVL2.5-8B) achieves 72.1%, surpassing InternVideo2.5-8B and VideoTree by 8.2% and 11.0%.
-On LongVideoBench, VideoChat-A1 (InternVideo2.5-8B) achieves 65.4%, surpassing ShareGPT4Video-8B and Gemini 1.5Pro by 25.7% and 1.4%.
-On MLVU, VideoChat-A1 (InternVideo2.5-8B) achieves 76.2%, outperforming GPT-4o by 11.6%.
-On VideoMME, VideoChat-A1 (InternVL2.5-8B) achieves 72.3% (w/o subs) and 77.0 (w/ subs), outperforming InternVL-2.5-78B by 0.2% and 3.0%.
-Efficiency Analysis.
-Table 2 presents a comparison of the average frame number and average inference time across various models.
-Compared to both closed-source models and large open-source models, VideoChat-A1 achieves substantial gains in efficiency.
-For instance, VideoChat-A1 (InternVL2.5-8B) processes only 35.9 frames and completes inference in 14.7 seconds on average, reducing frame usage by over 90% and inference time by more than 85% compared to GPT-4o, Gemini 1.5 Pro, and Qwen2.5-VL-72B.
+For example, on MLVU, VideoChat-A1 (InternVideo2.5-8B) achieves 76.2%, outperforming GPT-4o by 11.6%.
+Table 2 compares the average frame number and inference time.
+VideoChat-A1 achieves substantial gains in efficiency.
+For instance, VideoChat-A1 (InternVL2.5-8B) reduces frame usage by over 90% and inference time by more than 85% compared to GPT-4o, Gemini 1.5 Pro, and Qwen2.5-VL-72B.
 Despite these significant savings, our method maintains competitive performance across benchmarks.
-In contrast to small open-source models such as Qwen2.5-VL-7B, VideoChat-A1 not only delivers significantly better performance but also requires fewer video frames and achieves faster or comparable inference.
-4.3 Ablation Study
+In contrast to small open-source models, VideoChat-A1 not only delivers significantly better performance but also requires fewer video frames and achieves faster or comparable inference.
+Ablation Study
 7/8B Agent
@@
 70.1 65.4 76.2 72.9/76.8
-Table 3: Ablation on the effect of Video Glance in VideoChat-A1.
+Table 3: Ablation on the Video Glance in VideoChat-A1.
 Ablation study of Video Glance.
-In this study, we conducted an ablation study to verify the impact of the video glance on the performance of the model.
+In this study, we conducted an ablation study to verify the impact of the video glance.
 Table 3 shows that the model with the video glance achieves higher scores on all the tested tasks.
@@
 Ablation on Shot Partition and Chain-of-Shot.
-As shown in Table 5, we conducted an ablation study on VideoChat-A1 (Qwen2.5-VL-7B) to evaluate the impact of Shot Partition (SP) and Chain-of-Shot (CoS) across Egoschema, LongVideoBench, MLVU, and VideoMME datasets.
+In Table 4, we evaluate the impact of Shot Partition (SP) and Chain-of-Shot (CoS) on VideoChat-A1 (Qwen2.5-VL-7B).
 The EgoSchema and LongVideoBench are abbreviated as Ego and LVbench in the table because of the limited space.
-For a fair comparison, in the setup without shot partition, we evenly divided the entire video into six shots on average in the 1st round.
+In the setup without SP, we evenly divided the entire video into six shots on average in the 1st round.
 In subsequent rounds, for the selected shots that need to be expanded, they are directly partitioned into two shots on average.
-Regarding the setup without Chain-of-Shot, we removed the Chain-of-Shot mechanism.
-Instead, the model was made to answer three rounds.
-In each round, the model samples additional frames from the video and incorporates them into the existing frame set from the previous round.
-Then we conducted a vote, selecting the answer that appeared most frequently as the final response.
-Individually, enabling shot partition or Chain-of-Shot boosts performance.
+Regarding the setup without CoS, we removed the CoS mechanism.
+Instead, the model samples additional frames from the video and incorporates them into the existing frame set from the previous round.
+Then we conducted a vote to get the final response.
+Individually, enabling SP or CoS boosts performance.
 Their combined use yields the best results.
-For example, on MLVU, the score rises from 66.2 (with neither) to 71.9 (both enabled), highlighting their complementary roles in enhancing long-video understanding.
-SP
-CoS Ego LVBench MLVU VideoMME
+For example, on MLVU, the score rises from 66.2 (with neither) to 71.9 (both enabled).
+SP CoS Ego LVBench MLVU VideoMME
 67.1 56.4 66.2 63.4/69.8
@@
 70.7 64.2 71.9 71.8/76.5
-Table 4: Ablation of Shot Partition(SP) and Chain-of-Shot(CoS) in VideoChat-A1.
+Table 4: Ablation of Shot Partition (SP) and Chain-of-Shot (CoS) in VideoChat-A1 (Qwen2.5-VL-7B).
 Method of
@@
 Shot Partition 70.7 64.2 71.9 71.8/76.5
-Table 5: Ablation of different method of shot partition in VideoChat-A1 (Qwen2.5-VL-7B).
+Table 5: Ablation of different methods of shot partition in VideoChat-A1 (Qwen2.5-VL-7B).
 Max Rounds Ego LVBench MLVU VideoMME
@@
 9 69.3 64.0 71.8 71.5/76.2
-Table 7: Ablation over the number of initialized shots in VideoChat-A1 (Qwen2.5-VL-7B).
+Table 7: Ablation over number of initialized shots in VideoChat-A1 (Qwen2.5-VL-7B).
 Ablation on Shot Partition Methods.
 As shown in Table 5, we evaluate the impact of three methods of shot partition across multiple benchmarks.
-Cluster Partition refers to the clustering approach used in VideoTree [46], where all frames within each cluster are treated as a single shot, even though they may not form a continuous video segment.
+Cluster Partition refers to the clustering approach used in VideoTree (Wang et al.
+2024c), where all frames within each cluster are treated as a single shot, even though they may not form a continuous video segment.
 Average Partition refers to dividing each shot into smaller shots by uniformly splitting it into equal parts.
-Our proposed Shot Partition method consistently outperforms both Cluster Partition and Average Partition across all four benchmarks.
-These results demonstrate that partitioning videos into temporally coherent shots based on semantic information leads to significantly better understanding of long videos.
-Our Shot Partition achieves best accuracies of 70.7, 64.2, 71.9, and 71.8/76.5 on EgoSchema, LongVideoBench, MLVU, and VideoMME.
-Compared to the better of the other two methods, our Shot Partition method achieves gains of 1.6%, 5.8%, 4.2%, and 4.0%/3.7%
+Our proposed Shot Partition method consistently outperforms the other two methods, which demonstrate that partitioning videos into temporally coherent shots based on semantic information leads to better understanding of long videos.
 Ablation Max Rounds of Reflection.
-We evaluate the impact of varying the maximum number of reflection rounds across multiple benchmarks.
-As shown in Table 7, increasing the number of rounds consistently leads to performance improvements on all the tested dataset.
-For example, the accuracy on EgoSchema increases from 67.3% when using one reflection round to 70.7% with three rounds, while the accuracy on LongVideoBench improves from 60.1% to 64.2%.
+We evaluate the impact of varying the maximum number of reflection rounds.
+As shown in Table 6, increasing the number of rounds consistently leads to performance improvements.
 In the context of multi-round discussions, increasing the number of discussion rounds tends to enhance the effectiveness.
@@
 Ablation on Number of Initialized Shots.
-To evaluate the impact of the number of initial shots on the experimental results, the initial shot count is set to 3, 6, and 9, respectively.
-As shown in Table 7, the 6-shot configuration achieves best accuracies of 70.7, 64.2, 71.9, and 71.8/76.5 on EgoSchema, LongVideoBench, MLVU, and VideoMME, respectively.
-These correspond to gains of 1.6%, 5.8%, 4.2%, and 4.0%/3.7% over the better of the other two configurations, demonstrating that the 6-shot setting yields the best overall performance.
-5 Conclusion
-In this paper, we introduced VideoChat-A1, a novel agent-based framework for effectively addressing long video understanding tasks.
-Unlike existing approaches that often overlook the intrinsic shot-based structure of long videos, our method explicitly employs a chain-of-shot reasoning paradigm.
+We evaluate the effect of the number of initial shots, which is set to 3, 6, and 9.
+As shown in Table 7, the 6-shot achieves best accuracies compared with the other two configurations.
+For the subsequent rounds, considering that iterative process generates many subshots and aiming to balance computational efficiency, we fix clustering number for each subshot at $K=2$.
+Conclusion
+In this paper, we introduced VideoChat-A1, an agent-based framework for effectively addressing long video understanding tasks.
+Unlike existing methods that often overlook the intrinsic shot-based structure , our method explicitly employs a Chain-of-Shot reasoning paradigm.
 Specifically, VideoChat-A1 progressively identifies and refines relevant video segments through three key iterative steps: Shot Selection, Shot Partition, and Shot Reflection.
-This iterative interaction mechanism closely mimics human cognitive processes, allowing the agent to thoughtfully reason and accurately answer user questions from long videos.
-Extensive experimental results demonstrate the effectiveness and efficiency of our approach, highlighting the substantial benefits of our method in long video understanding tasks.
-Appendix A Prompts of VideoChat-A1.
-In this section, we introduce the prompts used in VideoChat-A1, as shown in Table LABEL:tab:prompts.
-These prompts are carefully designed to precisely fulfill the requirements of each step in the VideoChat-A1 pipeline.
-Table 8: The prompting templates used in different steps of VideoChat-A1.
-Prompt Content
-Prompt to decide whether to watch the whole video You are given a single-choice question, options, subtitles, and some frames of the long video.
-You should not only look at the textual information but also consider the input visual information, taking everything into account.
-If you can answer the question accurately and comprehensively based on the existing information, especially the visual information, and further watching the entire video will not significantly improve the quality of the answer, then you don’t need to watch the entire video and can answer ’No.’.
-However, if the existing information is not sufficient to fully answer the question, and watching the entire video may obtain information crucial for answering the question, please reply ’Yes’.
-The frame tokens: {Frame tokens}
-The question is:{Question}
-The options is:{Options}
-The subtitiles is:{Subtitiles}(If have)
-Output:{Yes/No}
-Prompt for generate the key information Given some frames from a long video, subtitles, a single-choice question, and options, identify the key information needed to answer the question.
-Focus on visual cues, context, and temporal relationships within the frames.
-Limit your response to 50 words.
-The frame tokens: {Frame tokens}
-The question is:{Question}
-The options is:{Options}
-The subtitiles is:{Subtitiles}(If have)
-Prompt for generating the answer Given some frames from a long video, subtitles, a single-choice question, and options, select the best answer to the following question based on the video and the subtitles.
-Respond with only the letter (A, B, C, or D) of the correct option.
-The frame tokens: {Frame tokens}
-The question is:{Question}
-The options is:{Options}
-The subtitiles is:{Subtitiles}(If have)
-The Key Info is:{Key Info}
-The best answer is:
-Prompt for generating the reason You are given a single-choice question, options, some frames of the long video, the key information of the question and the choice you have made.
-You should not only look at the textual information but also consider the input visual information, taking everything into account.
-Base on the provided information, your task is to explain the reason behind your choice.
-The frame tokens: {Frame tokens}
-The question is: {Question}
-The options is:{Options}
-The Key Info is:{Key Info}
-Your choice is {Choice}
-Prompt for confidence Evaluation You are given a single-choice question, options, some frames of the long video, your choice for the question and the reason.
-Your task is to evaluate whether the selected choice is correct.
-Criteria for Evaluation: Insufficient Information (Confidence Level: 1): There isn’t enough clear information to be sure.
-Partial Information (Confidence Level: 2): The decision is likely correct, but there’s some uncertainty.
-Sufficient Information (Confidence Level: 3): The decision is clearly supported by the given information.
-You need to provide the Confidence Level (a number of 1 or 2 or 3) according to the Criteria for Evaluation.
-Choose the confidence level that best reflects how clearly the input supports your choice.
-Respond with only the JSON format: ’confidence’: ’x’, where ’x’ is Confidence Level, without any additional text, explanation, or formatting.
-The frame tokens: {Frame tokens}
-The question is: {Question}
-The options is:{Options}
-Your choice is {Choice}
-Your reason is {Reason}
-Prompt for updating the key information in Round i Given some frames from a long video, subtitles, a single-choice question, and options, the old key information and the old choices and reasons, identify the key information needed to answer the question.
-Focus on visual cues, context, and temporal relationships within the frames.
-Limit your response to 50 words.
-The frame tokens: {Frame tokens}
-The question is: {Question}
-Round 1 Key Info: {Key Info in Round 1}
-Round 1 Choice: {Choice in Round 1}
-Round 1 Reason: {Reason in Round 1}
-Round i-1 Key Info: {Key Info in Round i-1}
-Round i-1 Choice: {Choice in Round i-1}
-Round i-1 Reason: {Reason in Round i-1}
-Appendix B Visualization and Experimental Comparison of Different Models.
-As shown in Fig.
-5, direct reasoning models such as GPT-4o [34] and Qwen2.5VL-72B [1] perform global sampling on videos and struggle to focus on key information in long videos.
-Agent-based methods like VideoTree [46] often suffer from incorrect or redundant sampling to generate noisy caption for wrong answer, due to the lack of deep thinking on shots in a long video.
-Although VideoAgent [43] implements a dynamic frame selection process, it performs retrieval in a frame-by-frame manner, which makes it susceptible to redundant information.
-Moreover, VideoAgent relies heavily on captions to answer questions, resulting in the loss of substantial visual information.
-While CoS [13] shares the same name—Chain of Shot—with the method proposed in this work, it does not implement a multi-round dynamic shot partition.
-Instead, it instead performs only a single-step inference.
-Specifically, it uniformly samples frames from the video and groups every four adjacent frames into a single “shot.” All shots are then categorized into positive and negative shots based on their relevance to the given question.
-The model generates three separate answers: one from the positive shots, one from the negative shots, and one from uniformly sampled frames.
-These responses are aggregated to produce the final answer.
-In contrast, VideoChat-A1 interactively employs shot perception and reasoning via Chain-of-Shot, which progressively looks into relevant shots through a reflective process to achieve superior performance.
-Table 9 presents the experimental results of our method compared with the agent-based methods shown in Fig.
-5.
-VideoChat-A1 (InternVL2.5-8B) achieves 72.1% on Egoschema, corresponding to absolute gains of 18.0% and 11.0% over VideoAgent (54.1%) and VideoTree (61.1%), respectively.
-On LongVideoBench, VideoChat-A1 (InternVideo2.5-8B) attains 65.4%, improving by 12.6% over CoS (LongVA) and by 6.5% over CoS (LLaVA-Video).
-For MLVU (M-avg), VideoChat-A1 (InternVideo2.5-8B) reaches 76.2%, outperforming CoS (LongVA) by 17.3% and CoS (LLaVA-Video) by 4.8%.
-Finally, on Video-MME without subtitles, VideoChat-A1 (InternVideo2.5-8B) achieves 72.9%, with 19.4% and 8.5% improvements over CoS (LongVA) and CoS (LLaVA-Video).
-On Video-MME with subtitles, it achieves 76.8%, with 21.1% and 5.3% improvements over CoS (LongVA) and CoS (LLaVA-Video), respectively.
-Model Egoschema LongVideoBench MLVU Video-MME
-(M-avg) Short Medium Long Average
-VideoAgent [43] 54.1 - - - - - -
-VideoTree (GPT-4)  [46] 61.1 - - - - - -
-CoS (LongVA[13] - 52.8 58.9 61.6/64.2 52.0/54.4 46.8/48.5 53.5/55.7
-CoS (LLaVA-Video[13] - 58.9 71.4 77.2/80.1 62.4/69.4 53.8/65.1 64.4/71.5
-VideoChat-A1 (Qwen2.5-VL-7B) 70.7 64.2 71.9 78.0/81.2 73.1/77.4 64.3/70.9 71.8/76.5
-VideoChat-A1 (InternVL2.5-8B) 72.1 65.2 75.1 80.8/81.4 72.3/76.8 63.7/72.6 72.3/77.0
-VideoChat-A1 (InternVideo2.5-8B) 70.1 65.4 76.2 81.4/82.4 72.8/76.7 65.0/71.2 72.9/76.8
-Table 9: Comparison with other three agent-baesd models on four Long Video Understanding Tasks. The Video-MME results are presented in the format “w/o subs / w/ subs".
-Figure 5: Visual comparison of different models.
-Appendix C Details of the Training of LongCLIP.
-To enable LongCLIP to handle long-video understanding tasks, we construct a fine-tuning dataset by sampling videos from ActivityNet [17], OpenVid-1M [33], ViTT [15], and MovieChat-Caption [40].
-Specifically, we first select videos that are already segmented into multiple clips, each paired with corresponding captions.
-These clip-caption pairs are directly used as training samples.
-For the remaining videos, we filter out those shorter than 5 seconds, longer than 15 minutes, or whose captions contain fewer than 10 words.
-After filtering, we manually create a subset, resulting in a long-video retrieval dataset consisting of 47k samples.
-The average video length is 101.40 seconds, and each caption contains an average of 72.29 words.
-The composition of the dataset is summarized in Table 10.
-Dataset Instance Num Avg Caption Length Avg Video Length
-ActivityNet-Caption 2,1352 15.20 165.32
-OpenVid-1M 2,2348 132.22 8.78
-ViTT 3,024 22.51 280.04
-MovieChat-Caption 496 132.92 432.57
-Total 47,220 72.29 101.40
-Table 10: Instance numbers of different datasets for training the LongCLIP.
-Since LongCLIP is originally designed for image inputs, we adapt it to video by uniformly sampling 16 frames from each shot and applying a 3D convolution to extract spatiotemporal features compatible with LongCLIP’s input format.
-We adopt the training hyperparameters following the LongCLIP paper, with a batch size of 1024, 3 training epochs, a weight decay of 1e-2, and a learning rate of 1e-4.
+This iterative interaction mechanism closely mimics human cognitive processes, allowing the agent to thoughtfully reason and accurately answer the questions.
+Experimental results demonstrate the effectiveness and efficiency of our method, highlighting the substantial benefits in long video understanding tasks.
+Acknowledgements
+Supported by Shanghai Artificial Intelligence Laboratory, and the National Key R&D Program of China(NO.2022ZD0160505).